5. Assume that X is following a normal distribution with mean y and variance o?. We want to test: Ho :u = 40 versus H:ll = 45what if X<42, what's the answer of these four question?
Answer #1

Sony X~ N(4202) se ll=40 we reject null a Mo: M= 40 vs M M = 45 Given that If x>42 then 4:45 otherwise u=40 critical region i@ the Racer & the test (6-8) - P[Rejecting Holteis false? H oe 5 P[Rejecting Holu, is folue) than any .=P [6> 42.1 Hos Ksus]@ Find n and a such that en and C 2-0.01 and 3-0.05 ) -) 2 =P[ ž>c/40) = o.ol pl X-My Gr 1 Mo: M:40] = 0.01 val() Vol o A >>to be foons 31 (1) , we have 6-45 - 164 C-VO9:33 DFS 233 C -45 *233 = -164C+164640 =) 3976 = 164*40 +45 233 > =166** 40 745*